# The Chained Difference-in-Differences

### v3

@@ 2.2.1 Rotating panel data
$\tilde{w}^{G}_{\tau\tau-1}(g,X)=\frac{E(G_{g}S_{\tau}S_{\tau-1})}{E[S_{\tau}S_{\tau-1}|X,G_{g}]E(G_{g})}w^{G}_{\tau\tau-1}(g,X),$ (17)
$\tilde{w}^{C}_{\tau\tau-1}(g,X)=\frac{E(G_{g}S_{\tau}S_{\tau-1})}{E[S_{\tau}S_{\tau-1}|X,C]E(G_{g})}w^{C}_{\tau\tau-1}(g,X).$ (18)
Corollary 1 extends the two-step estimator derived from Theorem 1 to account for sample selection based on observables. This can be achieved by reweighting each difference $Y_{t}-Y_{t-1}$ with the corresponding factor $\frac{E(G_{g}S_{\tau}S_{\tau-1})}{E[S_{\tau}S_{\tau-1}|X,a]E(G_{g})}$ for $a\in\{G_{2},...,C\}$ before applying the same chained DiD estimator. By doing so, we control for potential biases arising from observable characteristics influencing sample selection in addition to unobservable factors, such as unobservable individual heterogeneity. [^18] Remark that one can also opt for applying the stabilized weights $P(S_{\tau}S_{\tau-1}|X_{i},a_{i})^{-1}/n^{-1}\sum_{i=1}^{n}P(S_{\tau}S_{\tau-1}|X_{i},a_{i})^{-1}$ ex-ante.
@@
$\tilde{\tilde{w}}^{G}_{\tau\tau-1}=\frac{E[G_{g}S_{\tau}S_{\tau-1}]}{E[S_{\tau}|X,Y_{\tau-1},S_{\tau-1}=1,G]E[S_{\tau-1}|X,G]E[G_{g}]}w^{G}_{\tau\tau-1}(g,X),$ (21)
$\tilde{\tilde{w}}^{C}_{\tau\tau-1}=\frac{E[G_{g}S_{\tau}S_{\tau-1}]}{E[S_{\tau}|X,Y_{\tau-1},S_{\tau-1}=1,C]E[S_{\tau-1}|X,C]E[G_{g}]}w^{C}_{\tau\tau-1}(g,X).$ (22)
@@ A.2.1 Proofs for rotating panel setting
$=\sum_{\tau=g}^{t}E\left(\Delta ATT_{X}(g,\tau)|G_{g}=1,S_{\tau,\tau-1}=1\right)$
$=\sum_{\tau=g}^{t}E\left(A_{X}(g,\tau)-B_{X}(g,\tau)|G_{g}=1,S_{\tau,\tau-1}=1\right)$
@@
$E\left(A_{X}(g,\tau)|G_{g}=1,S_{\tau,\tau-1}=1\right)$ $=E\left(Y_{\tau}-Y_{\tau-1}|G_{g}=1,S_{\tau,\tau-1}=1\right)$ (33)
$=E_{M}\big{(}\left(Y_{\tau}-Y_{\tau-1}\right)\frac{G_{g}S_{\tau,\tau-1}}{E[G_{g}S_{\tau}S_{\tau-1}]}\big{)}$
@@
$E(B_{X}(g,\tau)|G_{g}=1,S_{\tau,\tau-1}=1)=E(E[Y_{\tau}-Y_{\tau-1}|X,C,S_{\tau,\tau-1}]|G_{g}=1,S_{\tau,\tau-1}=1)$ (34)
$=E\left(E[\frac{C}{1-P(G_{g}=1|X,G_{g}+C,S_{\tau,\tau-1})}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C=1,S_{\tau,\tau-1}]|G_{g},S_{\tau,\tau-1}\right)$
$=\frac{E\left(G_{g}S_{\tau,\tau-1}E[\frac{C}{1-P(G_{g}|X,G_{g}+C,S_{\tau,\tau-1})}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C,S_{\tau,\tau-1}]|G_{g}+C,S_{\tau,\tau-1}\right)}{{P(G_{g}=1|G_{g}+C,S_{\tau,\tau-1})}},$
@@
... $=\frac{E\left(G_{g}S_{\tau,\tau-1}E[\frac{C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C,S_{\tau,\tau-1}]|G_{g}+C,S_{\tau,\tau-1}\right)}{{P(G_{g}=1|G_{g}+C,S_{\tau,\tau-1})}}$ (35)
$=\frac{E\left(E[\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C,S_{\tau,\tau-1}]|G_{g}+C,S_{\tau,\tau-1}\right)}{{E(G_{g}|G_{g}+C,S_{\tau,\tau-1})}}$
$=\frac{E\left((G_{g}+C)E[\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C,S_{\tau,\tau-1}]|S_{\tau,\tau-1}\right)}{{E(G_{g}|G_{g}+C,S_{\tau,\tau-1})E((G_{g}+C)|S_{\tau,\tau-1})}}$
$=\frac{E\left((G_{g}+C)E[\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C,S_{\tau,\tau-1}]|S_{\tau,\tau-1}\right)}{{E(G_{g}|S_{\tau,\tau-1})}}$
$=\frac{E\left(E[(G_{g}+C)|X,S_{\tau,\tau-1}]E[\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C,S_{\tau,\tau-1}]|S_{\tau,\tau-1}\right)}{{E(G_{g}S_{\tau,\tau-1})}}$
$=\frac{E\left(E[\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,S_{\tau,\tau-1}]|S_{\tau,\tau-1}\right)}{{E(G_{g}|S_{\tau,\tau-1})}}$
$=\frac{E\left(\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|S_{\tau,\tau-1}\right)}{{E(G_{g}|S_{\tau,\tau-1})}}$
$=\frac{E_{M}\left(\frac{p_{g}(X)CS_{\tau,\tau-1}}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})\right)}{{E_{M}(G_{g}S_{\tau,\tau-1})}}$
@@
$E_{M}\left(\frac{p_{g}(X)CS_{\tau,\tau-1}}{1-p_{g}(X)}\right)$ $=E_{M}(\frac{E_{M}(G_{g}|X,G_{g}+C=1)C}{E_{M}(C|X,G_{g}+C=1)}S_{\tau,\tau-1})$ (36)
$=E_{M}(\frac{E_{M}(G_{g}|X)E_{M}(C|X)}{E_{M}(C|X)}S_{\tau,\tau-1})$
$=E_{M}(E_{M}(G_{g}|X)S_{\tau,\tau-1})$
$=E_{M}(G_{g}S_{\tau,\tau-1}).$
@@ A.2.6 General missing data patterns
Therefore, one can estimate an attrition model ex-ante for each treatment group and the control group to obtain $E[S_{\tau}S_{\tau-1}|X,G_{g}]$ and $E[S_{\tau}S_{\tau-1}|X,C]$, then reweight all first differences in outcome by $\frac{E(G_{g}S_{\tau}S_{\tau-1})}{E[S_{\tau}S_{\tau-1}|X,G_{g}]E(G_{g})}$ and $\frac{E(G_{g}S_{\tau}S_{\tau-1})}{E[S_{\tau}S_{\tau-1}|X,C]E(G_{g})}$ before applying our chained DiD estimator. Notice that, in general, the attrition model does not require to use the same explanatory variable $X$ than the propensity score $p_{g}(X)$. Remark, however, that we must be able to identify $E(G_{g})$ in the population.
@@
$\tilde{w}^{G}_{\tau\tau-1}(g,X)=\frac{E(G_{g}S_{\tau}S_{\tau-1})}{E[S_{\tau}S_{\tau-1}|X,G_{g}]E(G_{g})}w^{G}_{\tau\tau-1}(g,X),$ (55)
$\tilde{w}^{C}_{\tau\tau-1}(g,X)=\frac{E(G_{g}S_{\tau}S_{\tau-1})}{E[S_{\tau}S_{\tau-1}|X,C]E(G_{g})}w^{C}_{\tau\tau-1}(g,X).$ (56)
@@
$l_{t}(q_{a})=\frac{E(G_{g}S_{\tau}S_{\tau-1})}{q_{a}(X)E(G_{g})},$ (57)
with $q_{a}(X_{i})=E[S_{\tau}S_{\tau-1}|X_{i},G_{g}]=\Lambda(X_{i}^{\prime}\rho_{a}^{0})$ a parametric propensity scores with $\Lambda(\cdot)$ being a known function (logit or probit), that can be parametrically estimated by maximum likelihood. Remark that we could use other $X$’s here. We denote $\hat{q}_{a}(X_{i})=\Lambda(X_{i}^{\prime}\hat{\rho}_{g})$ where $\hat{\rho}_{a}$ are estimated by ML, $\dot{q}_{a}=\partial q_{a}(u)/\partial u$, and $\dot{q}_{a}(X)=\dot{q}_{a}(X_{i}^{\prime}\rho_{a}^{0})$. Under this assumption, the estimated parameter $\hat{\rho}_{a}$ is asymptotically linear.
@@
$N_{gt}^{C}=-\frac{E_{M}\left[\frac{p_{g}(X)C}{1-p_{g}(X)}X\frac{E(G_{g}S_{\tau}S_{\tau-1})}{q_{C}(X;\overline{\pi})^{2}E(G_{g})}\dot{q}_{C}(X;\overline{\rho})\left[(Y_{it}-Y_{it-1})-E_{M}\left[\tilde{w}^{C}_{it,t-1}(g,X)(Y_{it}-Y_{it-1})\right]\right]\right]}{E_{M}[\frac{p_{g}(X)C}{1-p_{g}(X)}]},$
$N_{gt}^{G}=-\frac{E_{M}\left[G_{g}S_{t,t-1}X\frac{E(G_{g}S_{\tau}S_{\tau-1})}{q_{C}(X;\overline{\pi})^{2}E(G_{g})}\dot{q}_{C}(X;\overline{\rho})\left[(Y_{it}-Y_{it-1})-E_{M}\left[\tilde{w}^{C}_{it,t-1}(g,X)(Y_{it}-Y_{it-1})\right]\right]\right]}{E_{M}[G_{g}S_{t,t-1}]},$
@@
$-\hat{E}_{M}\left[w_{t}(p_{g})X\frac{E(G_{g}S_{\tau}S_{\tau-1})}{q_{C}(X;\overline{\pi})^{2}E(G_{g})}\dot{q}_{C}(X;\overline{\rho})(Y_{t}-Y_{t-1})\right]^{\prime}\left(\hat{\rho}_{C}-\rho^{0}_{C}\right)+o_{p}(n^{-1/2}),$
@@
$-\hat{E}_{M}\left[w_{t}(p_{g})X\frac{E(G_{g}S_{\tau}S_{\tau-1})}{q_{C}(X;\overline{\pi})^{2}E(G_{g})}\dot{q}_{C}(X;\overline{\rho})\right]^{\prime}\left(\hat{\rho}_{C}-\rho^{0}_{C}\right)+o_{p}(n^{-1/2}),$

### v4

@@ 2.2.1 Rotating panel data
$\tilde{w}^{G}_{\tau\tau-1}(g,X)=\frac{E[S_{\tau}S_{\tau-1}|G_{g}]}{E[S_{\tau}S_{\tau-1}|X,G_{g}]}w^{G}_{\tau\tau-1}(g,X),$ (17)
$\tilde{w}^{C}_{\tau\tau-1}(g,X)=\frac{E[S_{\tau}S_{\tau-1}|C]}{E[S_{\tau}S_{\tau-1}|X,C]}w^{C}_{\tau\tau-1}(g,X).$ (18)
Corollary 1 extends the two-step estimator derived from Theorem 1 to account for sample selection based on observables. This can be achieved by reweighting each difference $Y_{t}-Y_{t-1}$ with the corresponding factor $\frac{E[S_{\tau}S_{\tau-1}|a]}{E[S_{\tau}S_{\tau-1}|X,a]}$ for $a\in\{G_{2},...,C\}$ before applying the same chained DiD estimator. By doing so, we control for potential biases arising from observable characteristics influencing sample selection in addition to unobservable factors, such as unobservable individual heterogeneity. [^18] Remark that one can also opt for applying the stabilized weights $P(S_{\tau}S_{\tau-1}|X_{i},a_{i})^{-1}/n^{-1}\sum_{i=1}^{n}P(S_{\tau}S_{\tau-1}|X_{i},a_{i})^{-1}$ ex-ante.
@@
$\tilde{\tilde{w}}^{G}_{\tau\tau-1}=\frac{E[S_{\tau}S_{\tau-1}|G_{g}]}{E[S_{\tau}|X,Y_{\tau-1},S_{\tau-1}=1,G]E[S_{\tau-1}|X,G]}w^{G}_{\tau\tau-1}(g,X),$ (21)
$\tilde{\tilde{w}}^{C}_{\tau\tau-1}=\frac{E[S_{\tau}S_{\tau-1}|C]}{E[S_{\tau}|X,Y_{\tau-1},S_{\tau-1}=1,C]E[S_{\tau-1}|X,C]}w^{C}_{\tau\tau-1}(g,X).$ (22)
@@ 5 Conclusion
Acknowledgments
We are indebted to Serena Ng for her guidance, an associate editor, and two anonymous referees for their valuable comments and suggestions that greatly improved the quality of this work. We would also like to thank Brantly Callaway, Joel Cuerrier, Laurent Davezies, Clément de Chaisemartin, Xavier D’Haultefœuille, Yannick Guyonvarch, Xavier Jaravel, and all participants to seminars and conferences for insightful discussions and comments.
The R package cdid is available on CRAN.
Funding
This work was supported by the Social Sciences and Humanities Research Council of Canada (430-2022-00544).
@@ A.2.1 Proofs for rotating panel setting
$=\sum_{\tau=g}^{t}E\left(\Delta ATT_{X}(g,\tau)|G_{g}=1\right)$
$=\sum_{\tau=g}^{t}E\left(A_{X}(g,\tau)-B_{X}(g,\tau)|G_{g}=1\right)$
@@
$=E\left(E[Y_{\tau}-Y_{\tau-1}|X,G_{g}]|G_{g}\right)$
$=E\left(E[Y_{\tau}-Y_{\tau-1}|X,G_{g},S_{\tau}S_{\tau-1}]|G_{g}\right)$
$=E\left(E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|G_{g}]}(Y_{\tau}-Y_{\tau-1})|X,G_{g}]|G_{g}\right)$
$=E_{M}\big{(}\left(Y_{\tau}-Y_{\tau-1}\right)\frac{G_{g}S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|G_{g}]E[G_{g}]}\big{)}$
@@
$=E(E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}(Y_{\tau}-Y_{\tau-1})|X,C]|G_{g})$
$=E\left(E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}\frac{C}{1-P(G_{g}=1|X,G_{g}+C)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C=1]|G_{g}\right)$
$=\frac{E\left(G_{g}E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}\frac{C}{1-P(G_{g}|X,G_{g}+C)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C]|G_{g}+C\right)}{{P(G_{g}=1|G_{g}+C)}},$
@@
... $=\frac{E\left(G_{g}E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}\frac{C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C]|G_{g}+C\right)}{{P(G_{g}=1|G_{g}+C)}},$ (35)
$=\frac{E\left(E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C]|G_{g}+C\right)}{{E(G_{g}|G_{g}+C)}},$
$=\frac{E\left((G_{g}+C)E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C]\right)}{{E(G_{g}|G_{g}+C)E(G_{g}+C)}},$
$=\frac{E\left((G_{g}+C)E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C]\right)}{{E(G_{g})}},$
$=\frac{E\left(E[(G_{g}+C)|X]E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X,G_{g}+C]\right)}{{E(G_{g})}}$
$=\frac{E\left(E[\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})|X]\right)}{{E(G_{g})}}$
$=\frac{E\left(\frac{S_{\tau}S_{\tau-1}}{E[S_{\tau}S_{\tau-1}|C]}\frac{p_{g}(X)C}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})\right)}{{E(G_{g})}}$
$=\frac{E\left(\frac{E(G_{g}S_{\tau}S_{\tau-1})}{E[S_{\tau}S_{\tau-1}|C]E(G_{g})}\frac{p_{g}(X)CS_{\tau}S_{\tau-1}}{1-p_{g}(X)}(Y_{\tau}-Y_{\tau-1})\right)}{{E(G_{g}S_{\tau}S_{\tau-1})}}$
@@
$E_{M}\left(\frac{p_{g}(X)CS_{\tau,\tau-1}}{1-p_{g}(X)}\right)$ $=E_{M}(\frac{E_{M}(G_{g}|X,G_{g}+C=1)}{E_{M}(C|X,G_{g}+C=1)}CS_{\tau,\tau-1})$ (36)
$=E_{M}(\frac{E_{M}(G_{g}|X)E_{M}(CS_{\tau,\tau-1}|X)}{E_{M}(C|X)})$
$=E_{M}(\frac{E_{M}(G_{g}|X)E(S_{\tau,\tau-1}|C,X)E_{M}(C|X)}{E_{M}(C|X)})$
$=E_{M}(E_{M}(G_{g}|X)E(S_{\tau,\tau-1}|C,X))$
$=E(S_{\tau,\tau-1}|C)E_{M}(G_{g}).$
@@ A.2.6 General missing data patterns
Therefore, one can estimate an attrition model ex-ante for each treatment group and the control group to obtain $E[S_{\tau}S_{\tau-1}|X,G_{g}]$ and $E[S_{\tau}S_{\tau-1}|X,C]$, then reweight all first differences in outcome by $\frac{E(S_{\tau}S_{\tau-1}|C)}{E[S_{\tau}S_{\tau-1}|X,C]}$ and $\frac{E(S_{\tau}S_{\tau-1}|G_{g})}{E[S_{\tau}S_{\tau-1}|X,G_{g}]}$ before applying our chained DiD estimator. Notice that, in general, the attrition model does not require to use the same explanatory variable $X$ than the propensity score $p_{g}(X)$. Remark, however, that we must be able to identify $E(G_{g})$ in the population.
@@
$\tilde{w}^{G}_{\tau\tau-1}(g,X)=\frac{E(S_{\tau}S_{\tau-1}|G_{g})}{E[S_{\tau}S_{\tau-1}|X,G_{g}]}w^{G}_{\tau\tau-1}(g,X),$ (55)
$\tilde{w}^{C}_{\tau\tau-1}(g,X)=\frac{E(S_{\tau}S_{\tau-1}|C)}{E[S_{\tau}S_{\tau-1}|X,C]}w^{C}_{\tau\tau-1}(g,X).$ (56)
@@
$l_{t}(q_{a})=\frac{E(S_{\tau}S_{\tau-1}|a)}{q_{a}(X)},$ (57)
with $q_{a}(X_{i})=E[S_{\tau}S_{\tau-1}|X_{i},a]=\Lambda(X_{i}^{\prime}\rho_{a}^{0})$ a parametric propensity scores with $\Lambda(\cdot)$ being a known function (logit or probit), that can be parametrically estimated by maximum likelihood. Remark that we could use other $X$’s here. We denote $\hat{q}_{a}(X_{i})=\Lambda(X_{i}^{\prime}\hat{\rho}_{a})$ where $\hat{\rho}_{a}$ are estimated by ML, $\dot{q}_{a}=\partial q_{a}(u)/\partial u$, and $\dot{q}_{a}(X)=\dot{q}_{a}(X_{i}^{\prime}\rho_{a}^{0})$. Under this assumption, the estimated parameter $\hat{\rho}_{a}$ is asymptotically linear.
@@
$N_{gt}^{C}=-\frac{E_{M}\left[\frac{p_{g}(X)C}{1-p_{g}(X)}X\frac{E(S_{\tau}S_{\tau-1}|C)}{q_{C}(X;\overline{\pi})^{2}}\dot{q}_{C}(X;\overline{\rho})\left[(Y_{it}-Y_{it-1})-E_{M}\left[\tilde{w}^{C}_{it,t-1}(g,X)(Y_{it}-Y_{it-1})\right]\right]\right]}{E_{M}[\frac{p_{g}(X)C}{1-p_{g}(X)}]},$
$N_{gt}^{G}=-\frac{E_{M}\left[G_{g}S_{t,t-1}X\frac{E(S_{\tau}S_{\tau-1}|G_{g})}{q_{G}(X;\overline{\pi})^{2}}\dot{q}_{G}(X;\overline{\rho})\left[(Y_{it}-Y_{it-1})-E_{M}\left[\tilde{w}^{C}_{it,t-1}(g,X)(Y_{it}-Y_{it-1})\right]\right]\right]}{E_{M}[G_{g}S_{t,t-1}]},$
@@
$-\hat{E}_{M}\left[w_{t}(p_{g})X\frac{E(S_{\tau}S_{\tau-1}|C)}{q_{C}(X;\overline{\pi})^{2}}\dot{q}_{C}(X;\overline{\rho})(Y_{t}-Y_{t-1})\right]^{\prime}\left(\hat{\rho}_{C}-\rho^{0}_{C}\right)+o_{p}(n^{-1/2}),$
@@
$-\hat{E}_{M}\left[w_{t}(p_{g})X\frac{E(S_{\tau}S_{\tau-1}|C)}{q_{C}(X;\overline{\pi})^{2}}\dot{q}_{C}(X;\overline{\rho})\right]^{\prime}\left(\hat{\rho}_{C}-\rho^{0}_{C}\right)+o_{p}(n^{-1/2}),$